### v1

# Clickbait Detection via Large Language Models

Abstract
Clickbait, which aims to induce users with some surprising and even thrilling headlines for increasing click-through rates, permeates almost all online content publishers, such as news portals and social media. Recently, Large Language Models (LLMs) have emerged as a powerful instrument and achieved tremendous success in a serious of NLP downstream tasks. However, it is not yet known whether LLMs can be served as a high-quality clickbait detection system. In this paper, we analyze the performance of LLMs in the few-shot scenarios on a number of English and Chinese benchmark datasets. Experimental results show that LLMs cannot achieve the best results compared to the state-of-the-art deep and fine-tuning PLMs methods. Different from the human intuition, the experiments demonstrated that LLMs cannot make satisfied clickbait detection just by the headlines.
1 Introduction
With the rapid development of online applications, some content publishers try to utilize clickbait for generating profits Chen et al. (2015). Clickbait refers to deliberately enticed users to click with some curious and chilling headlines, which are always unrelated to the real content or even the advertising promotion Chakraborty et al. (2016). The popularity of clickbait will inevitably lead to the experience degradation or even the disgust of users, there is an urgent demand to develop effective automatic clickbait detection methods Liu et al. (2022).
In recent years, the research methods for clickbait detection evolved from feature engineering to neural networks and, more recently, into pre-trained language models. Feature engineering methods extracted features such as semantic and linguistic features for detection tasks Biyani et al. (2016); Wei and Wan (2017). The deep neural networks methods can learn more abstract and higher-level features by disentangling explanatory factors of variations behind news title and content for clickbait detection Yoon et al. (2019); Zheng et al. (2021). In recent years, Pre-trained Language Models (PLMs), such as BERT Devlin et al. (2018), also show superiority in clickbait detection task. However, the feature engineering-based methods and deep neural networks typically require large-scaled labeled data since the detection is regarded as a classification method. In the PLMs methods, the huge gap between pre-training and fine-tuning prevent detection tasks from fully utilizing pre-training knowledge.
More recently, Large Language Models (LLMs) have demonstrated the powerful ability in various NLP downstream tasks Brown et al. (2020); Chowdhery et al. (2022); Thoppilan et al. (2022), which can achieve awesome performance even in the few-shot and zero-shot scenarios. Nevertheless, it remains unclear how LLMs perform in clickbait detection tasks compared to current methods. To address this issue, in this paper, we conduct a systematic evaluation of the few-shot and zero-shot learning capability of LLMs. The experiments are validated on both English and Chinese open datasets, and we conducted an empirical comparison of the performance between ChatGPT with the GPT3.5 model (gpt-3.5-turbo) and other state-of-the-art methods such as prompt-tuning.
To the best of our knowledge, this is the first attempt to validate the performance of LLMs on clickbait detection, the works in this paper provide a preliminary evaluation including detection results and robustness. The key findings and insights are summarized as follows:
  • ChatGPT achieves unsatisfying results in the few-shot scenarios compared to the state-of-the-art clickbait detection methods. We found that fine-tuning PLMs and prompt-tuning can achieve better results just based on the titles, which is consistent with humans in the real-world.
  • ChatGPT is a monolithic model capable of supporting multiple languages, which makes it a comprehensive multilingual clickbait detection technique. After evaluating the performance of ChatGPT on the task of clickbait detection across two languages (English and Chinese), we observed that it achieved stable results on almost all evaluation metrics. This also confirms that LLMs can be adapted to other languages.
It is worth mentioning that the source code and all the results of this paper is available on https://github.com/zhuyiYZU/chatGPTforClickbait.
2 Related Work
2.1 Clickbait Detection
Clickbait detection is an emerging field which is attracting increasing attention in recent years. On most online services, such as e-commerce, social media, and news portals, more clicks means more profit and commercial revenue. Early clickbait detection methods mainly focus on extracting a variety of features for detection tasks, such as semantic Rony et al. (2017), linguistics Blom and Hansen (2015), and multi-modal features Chen et al. (2015). However, these methods require expert knowledge for feature selection, and the handcrafted features are limited in representing more abstract and higher-level information.
In recent years, deep neural network models facilitate crossing and combination among even more diverse and sophisticated features, which has shown fairly good performance in clickbait detection. The popular deep neural networks, such as Recurrent Neural Networks (RNN) Anand et al. (2017), Convolutional Neural Networks (CNN) Agrawal (2016), Attention Mechanism Mishra et al. (2020), and Graph Attention Networks Liu et al. (2022), have already been devoted to the clickbait detection tasks. Despite the success of deep learning methods, due to the requirements on large-scale labeled training datasets, lead to high costs in collecting eligible training data. Recently, some domain adaptation López-Sánchez et al. (2018) and data augmentation Yoon et al. (2019) methods have been proposed to address the issues, however, these methods may bring additional noise in detection tasks.
More recently, pre-trained language models (PLMs) such as BERT Devlin et al. (2018), RoBERTa Liu et al. (2019), and T5 Raffel et al. (2020) have emerged as a powerful instruments for language understanding and generation. Through the fine-tuning PLMs on the special downstream task, the rich knowledge distributed in PLMs can be stimulated to better serve downstream tasks including clickbait detection Indurthi et al. (2020); Yi et al. (2022). Despite the success of fine-tuning PLMs, some recent studies find one of its critical challenges is the significant gap of objective forms in pre-training and fine-tuning, which restricts taking full advantage of knowledge in PLMs.
2.2 Large Language Models
Represented by GPT-3 Brown et al. (2020), Large Language Models (LLMs) have achieved superior performance, especially in few-shot learning scenarios Brown et al. (2020); Chowdhery et al. (2022); Thoppilan et al. (2022). Different from the previous PLMs methods, LLMs has two distinct advantages. The first is the larger scale, LLMs have a much larger scale in terms of model parameters and training data. Secondly, without the fine-tuning PLMs, LLMs can prompt few-shot learning that requires no additional neural layer and shows excellent performance. However, there is no work on the capabilities of LLMs on clickbait detection tasks.
3 Methodology
Through the few-shot learning ability of LLMs, they can also achieve excellent performance for tasks with low resources. Considering the lack of large-scale training corpus for clickbait detection tasks, we will test the performance of LLMs in few-shot scenarios on clickbait detection.
LLMs typically use prompts (i.e. specific templates) to guide the model in predicting output or answers, without requiring specific training on the data. Utilizing this form of prompt, we conducted experiments with different prompts on OpenAI’s largest available model GPT3.5 (gpt-3.5-turbo) for ChatGPT.
Prompt. To stimulate the rich knowledge distributed in LLMs, we manually designed the prompts for validating the performance of clickbait detection. The details of the prompts are illustrated in Table 1. For the two prompts, {Clickbait Sentence} and {Not-Clickbait Sentence} refer to the input sentence, {Yes, it is a clickbait} and {No, it is not a clickbait} are the corresponding labels respectively. {Results} is the place the carries the outputs of LLMs.
Table 1: The hand-crafted prompts for clickbait detection.
Prompts
P1
I want you to detect whether
the input sentence is clickbait or not.
Input: {Clickbait Sentence}
Output: {Yes, it is a clickbait}
Input: {Not-Clickbait Sentence}
Output: {No, it is not a clickbait}
Input: {Clickbait or Not-Clickbait Sentence}
Output: {Results}
P2 Sentence:{Clickbait Sentence}
Question: Detect the above sentence is clickbait or not.
Answer: {Yes, it is a clickbait}
Sentence:{Not-Clickbait Sentence}
Question: Detect the above sentence is clickbait or not.
Answer: {No, it is not a clickbait}
Sentence:{Clickbait or Not-Clickbait Sentence}
Question: Detect the above sentence is clickbait or not.
Answer: {Results}
In the first prompt (P1), the {guide-input-output} pattern is employed to guide LLMs for detecting clickbaits. In the second prompt (P2), the {sentence-question-answer} pattern is utilized to detect clickbaits in the form of a question. It is worth mentioning that the specialized guide word, such as ”Output:” and ”Answer:”, are added to the prompts, then the clickbait detection is regarded as the close-style task for LLMs, which ensured to achievement of a unique output. When performing multilingual clickbait detection tasks, we translate these two prompts into Chinese used in specific tasks as shown in Table 2.
Table 2: The hand-crafted prompts for clickbait detection.
提示
P1
请帮我检测
以下输入是或不是点击诱饵。
输入: {点击诱饵句子}
输出:{是,这是一个点击诱饵}
输入: {非点击诱饵句子}
输出: {不,这不是一个点击诱饵}
输入: {点击诱饵或非点击诱饵句子}
输出: {结果}
P2 句子:{点击诱饵句子}
问题: 检测上面的句子是或不是点击诱饵。
回答: {是,这是一个点击诱饵}
句子:{非点击诱饵句子}
问题: 检测上面的句子是或不是点击诱饵。
回答: {不,这不是一个点击诱饵}
句子:{点击诱饵或非点击诱饵句子}
问题: 检测上面的句子是或不是点击诱饵。
回答: {结果}
Zero-shot. In the clickbait detection experiments in zero-shot scenarios, just one {guide-input-output} or {sentence-question-answer} pattern is directly input into the LLMs for the detection result.
Few-shot. In the few-shot scenario, some instances follow the {guide-input-output} or {sentence-question-answer} pattern are provided as the training data for the detection tasks. Notably, in the first prompt (P1), the guidance is not needed to be repeated, only the input-output is repeated for some instance sentences and corresponding labels are stacked. In the second prompt (P2), the sentence-question-answer is repeated for each sample as training data.
(a) English P1
(b) Chinese P1
(c) English P2
(d) Chinese P2
Figure 1: The example of ChatGPT detecting clickbait errors.
4 Experiment
4.1 Datasets and Templates
To evaluate the performance of our method for clickbait detection in both English and Chinese, we conduct experiments on seven datasets. The DL-Clickbait (DLC), Clickbait news detection (CND), SC-Clickbait (SCC) are three well-known public clickbait detection datasets in English. Sina, Tencent, Wechat, and Paper are four public clickbait detection datasets in Chinese. The statistical details of all the datasets are presented in Table 3.
Table 3: Statistics of the news datasets. ”#” and ”avg.#” denote ”the number of” and ”the average number of”.
Statistic DLC CND SCC Sina Tencent Wechat Paper
#total headline 2345 8324 32000 1912 2546 2800 30800
#train sets 1638 5801 22400 1338 1782 1960 21560
#test sets 707 2523 9600 574 764 840 9240
avg.# words per sentence 11 13 9 21 25 21 22
4.2 Baselines
We compare ChatGPT with the following deep neural network and fine-tuning PLMs methods.
TextCNN Agrawal (2016) TextCNN utilizes kernels with different sizes to find various features from the input text. These features are automatically detected and are used to train a neural network for the corresponding task. It understands the input text from different perspectives, and can be applied to various nature language processing tasks.
BiLSTM Kumar et al. (2018) A bidirectional LSTM with an attention mechanism to learn the extent to which a word contributes to the clickbait score of a social media post in a different way. It used a Siamese net to detect similarities between the source and target data. To add another layer of complexity to the model, it also use TextCNN to learn image embeddings from large amounts of data.
Gated_CNN Yang et al. (2019) The method can be stacked to represent large context sizes and extract hierarchical features over larger and larger contexts with more abstractive features.
FAST_TEXT Joulin et al. (2016) FAST_TEXT is a text-classification method, and the sequence of words is considered in it, which is a model learning distributed representations of words based on ordered words.
BERT Devlin et al. (2018) The BERT method uses words and sentences to distinguish the context of words in a sentence.
4.3 Implementation details and evaluation metrics
Few-shot settings
In this experiment, we randomly selected K(5,10,20) instances as the training set for ChatGPT, and part of the samples are selected as the test set. Considering that too few training samples may greatly affect the effectiveness of the baselines, we draw different training samples from different datasets, where the number of training samples is 200/300/600, 300/600/1200, and 500/1000/2000 for English datasets (DL-Clickbait, CND, and SC-Clickbait datasets) and 200/400/800, 400/800/1600, 400/800/1600 and 800/1600/3200 for Chinese datasets (Sina, Tencent, Wechat and Paper), corresponding to 5/10/20 shots in ChatGPT. Considering that different choices of few-shot training during training affect the test results, we repeatedly sample the same data on K random seeds simultaneously and calculate their mean values as reported. Considering more intuitive approaches to training the ChatGPT model, here are a few specific examples. As shown in Figure 1, examples are provided for training the model using few-shot (5-shot) methods, along with the corresponding results.
Table 4: The Accuracy, Precision, Recall, F1-scores results (%) on both English (DLC, CND and SCC) and Chinese (Sina, Tencent, Wechat and Paper) clickbait datasets in zero-shot scenario.
metrics Dataset
CND SCC DLC Sina Tencent Wechat Paper
Acc 63.66 53.87 76.44 61.23 53.80 62.02 63.59
Pre 47.88 53.69 75.87 54.61 53.13 57.71 61.71
Rec 74.79 49.68 77.53 57.84 64.40 90.00 71.54
F1 58.38 51.61 76.69 57.18 58.22 70.33 66.26
Table 5: The Accuracy, Precision, Recall, F1-scores results (%) on English datasets (DLC, CND and SCC). 200/5,300/10,and 600/20 denote the training number of not-prompt-tuning(CNN, BILSTM, GATED_CNN, FAST_TEXT and BERT) and ChatGPT methods. Ditto above.
Data Method Shot
200/5 300/10 600/20
Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
CND CNN 77.86 79.36 77.77 79.27 80.50 79.00 80.83 79.12 80.59 80.70 80.64 80.35
BiLSTM 77.28 78.71 79.83 75.49 79.11 79.07 78.37 78.72 79.10 82.19 80.16 80.49
GATED_CNN 77.80 80.05 75.41 78.01 79.34 79.86 78.16 78.43 78.39 79.71 79.93 79.66
FAST_TEXT 65.23 72.31 63.84 72.32 74.89 71.07 74.30 65.58 75.13 78.33 73.60 77.63
BERT 71.25 72.00 71.25 71.00 77.54 78.28 76.46 77.49 82.52 83.16 81.08 82.94
ChatGPT 80.75 69.06 82.79 73.56 79.33 69.51 88.02 70.88 82.07 75.66 82.64 73.67
300/5 600/10 1200/20
Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
SCC CNN 73.74 69.70 71.57 68.82 76.47 74.83 74.33 75.70 73.19 72.05 75.27 66.61
BiLSTM 75.18 75.78 74.06 73.50 75.74 79.54 78.28 79.15 71.79 81.70 80.08 82.71
GATED_CNN 70.92 71.41 70.18 67.69 71.26 74.53 76.04 72.28 68.00 74.54 74.23 75.39
FAST_TEXT 54.18 66.01 64.12 61.22 72.51 71.61 69.57 53.51 79.25 77.76 74.47 79.69
BERT 73.17 73.19 73.17 73.16 76.90 80.65 76.90 76.71 83.28 85.69 83.02 83.80
ChatGPT 60.00 59.02 68.36 62.06 59.44 60.76 61.00 60.13 60.90 63.33 66.88 62.57
500/5 1000/10 2000/20
Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
DLC CNN 91.59 91.51 91.56 91.66 93.32 93.04 93.19 92.84 94.53 94.14 94.71 94.72
BiLSTM 86.69 86.45 85.91 86.48 87.36 86.08 88.20 87.97 89.58 88.45 88.24 89.14
GATED_CNN 92.07 91.78 91.90 91.67 93.56 93.14 93.27 93.56 93.22 94.67 94.28 94.45
FAST_TEXT 84.09 83.66 83.18 82.52 85.99 85.76 85.27 85.74 86.67 86.96 86.60 86.68
BERT 78.82 79.23 78.16 78.18 85.81 88.91 85.81 85.83 91.42 92.54 91.42 91.37
ChatGPT 83.94 88.63 90.01 83.08 86.14 92.41 86.93 85.09 88.73 95.54 87.81 87.82
Table 6: The Accuracy, Precision, Recall, F1-scores results (%) on Chinese datasets (Sina, Tencent, Wechat and Paper). 200/5,400/10,and 800/20 denote the training number of not-prompt-tuning(CNN, BILSTM, GATED_CNN, FAST_TEXT and BERT) and ChatGPT methods. Ditto above.
Data Method Shot
200/5 400/10 800/20
Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
Sina CNN 72.13 72.02 72.47 72.70 77.00 74.85 75.61 75.84 75.96 77.60 77.35 78.19
BiLSTM 74.74 71.83 58.71 70.88 74.39 72.32 72.65 75.40 75.60 76.11 72.65 74.01
GATED_CNN 73.87 72.09 71.25 71.37 75.61 76.07 77.35 75.66 77.00 76.20 75.61 76.93
FAST_TEXT 56.10 62.23 52.79 63.94 74.04 71.45 72.13 72.74 73.17 73.76 74.04 71.81
BERT 52.36 52.92 44.25 48.20 55.25 57.69 46.51 51.50 65.96 71.43 55.56 62.50
ChatGPT 70.73 76.21 60.28 67.32 69.16 71.65 63.41 67.28 75.30 83.80 62.85 71.83
400/5 800/10 1600/20
Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
Tencent CNN 54.58 53.97 56.94 54.42 59.55 56.16 59.16 57.41 61.65 59.29 59.69 62.75
BiLSTM 53.80 54.06 50.65 52.87 53.27 58.43 59.69 54.90 59.03 59.87 57.85 57.56
GATED_CNN 55.37 56.28 56.54 55.63 60.60 60.29 57.33 55.71 60.08 60.74 57.98 60.31
FAST_TEXT 53.53 50.58 51.05 56.49 56.28 57.75 57.07 49.95 55.37 57.94 58.12 58.20
BERT 47.71 47.55 43.19 45.27 52.80 54.20 47.26 50.49 59.54 62.00 53.26 57.30
ChatGPT 59.34 57.91 71.13 63.84 56.02 54.71 69.90 61.38 59.08 57.17 71.99 63.73
400/5 800/10 1600/20
Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
Wechat CNN 72.14 70.64 69.40 70.60 86.67 84.38 85.12 83.59 85.83 86.70 87.38 86.72
BiLSTM 64.29 61.93 67.74 62.24 81.67 81.52 76.55 81.61 83.57 81.78 80.95 83.49
GATED_CNN 67.74 69.17 69.17 67.62 83.45 86.31 84.29 85.27 87.74 88.16 86.67 86.02
FAST_TEXT 58.10 62.59 60.48 61.56 81.07 83.26 81.79 83.13 84.17 84.60 84.29 81.18
BERT 45.05 44.76 41.67 43.16 49.54 50.73 44.73 47.54 57.06 59.39 51.48 55.15
ChatGPT 60.12 56.69 85.71 68.24 60.41 56.16 93.99 70.26 64.48 59.93 90.24 71.01
800/5 1600/10 3200/20
Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
Paper CNN 68.87 73.27 68.76 66.11 71.45 74.28 70.47 71.16 72.38 72.12 71.32 72.29
BiLSTM 64.36 68.41 65.96 65.89 68.44 68.07 66.26 64.94 69.78 68.23 69.03 69.85
GATED_CNN 68.44 72.40 68.63 67.29 71.41 74.78 69.36 65.84 71.79 72.02 71.04 72.91
FAST_TEXT 67.93 70.25 66.03 65.53 68.45 72.88 68.96 67.85 71.10 72.15 72.23 70.39
BERT 50.16 50.20 40.02 44.54 51.92 52.21 41.51 46.26 54.17 54.48 43.60 48.44
ChatGPT 68.33 72.18 62.70 66.42 68.32 70.40 63.88 66.98 70.94 72.14 67.21 69.59
Implementation details.
We chose the latest available LLMs GPT3.5 (gpt-3.5-turbo). In order to interact with the model, we access it by making API calls to the ChatGPT API, in which calling GPT3.5 requires payment. This model’s maximum context length is 4096 tokens, approximating 100 news headlines. Moreover, this model can only call three news headlines a minute. In addition, for the few-shot experiments, we select a few sentences that ChatGPT got wrong from the sevendata sets mentioned above as examples of clickbait detection. Figure 1 shows ChatGPT detecting clickbait errors.
Table 7: Comparison of different prompts for GPT3.5 on English datasets (%).
Data Method Shot
5 10 20
Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
CND GPT3.5+P1 80.75 69.06 78.69 73.56 79.33 69.51 70.08 69.79 82.07 75.66 70.08 72.76
GPT3.5+P2 67.98 52.20 82.79 64.03 75.25 59.33 88.02 70.88 79.80 66.45 82.64 73.67
SCC GPT3.5+P1 58.41 59.02 49.36 53.76 59.44 60.76 50.08 54.91 60.90 63.33 49.20 55.38
GPT3.5+P2 60.00 56.83 68.36 62.06 59.00 59.28 61.00 60.13 60.63 58.78 66.88 62.57
DLC GPT3.5+P1 83.94 88.63 78.19 83.08 86.14 92.41 78.85 85.09 88.73 95.54 81.26 87.82
GPT3.5+P2 81.49 76.54 90.01 82.73 84.39 82.71 86.93 84.77 87.52 87.30 87.81 87.55
Table 8: Comparison of different prompts for GPT3.5 on Chinses datasets (%).
Data Method Shot
5 10 20
Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
Sina GPT3.5+P1 65.33 68.80 56.10 61.80 60.80 60.69 61.32 61.00 63.94 64.60 61.67 63.10
GPT3.5+P2 70.73 76.21 60.28 67.32 69.16 71.65 63.41 67.28 75.30 83.80 62.85 0.7183
Tencent GPT3.5+P1 59.34 57.91 71.13 63.84 52.88 52.89 52.62 52.75 58.51 57.17 67.80 62.03
GPT3.5+P2 58.64 56.96 70.68 63.08 56.02 54.71 69.99 61.38 59.08 57.17 71.99 63.73
Wechat GPT3.5+P1 60.12 56.69 85.71 68.24 60.41 56.10 93.99 70.26 58.57 55.25 90.24 68.54
GPT3.5+P2 55.71 53.99 77.38 63.60 58.69 56.16 79.29 65.75 64.48 59.93 87.11 71.01
Paper GPT3.5+P1 68.33 70.60 62.70 66.42 68.32 70.40 63.88 66.98 68.70 70.84 61.94 66.09
GPT3.5+P2 68.13 72.18 59.59 65.28 66.43 69.54 55.77 61.9 70.94 72.14 67.21 69.59
Evaluation metrics
To test the effect of detection, we conduct four evaluation metrics to evaluate our method in experiments, such as accuracy, precision, recall and F1-score.
accuracy The accuracy can be defined as the ratio of correctly predicted samples to the total number of samples.
$Acc=\frac{TP+TN}{TP+TN+FP+FN}$
precision The positive prediction rate can be defined as the ratio of correctly predicted positive samples to the total number of positive samples in the prediction.
$P=\frac{TP}{TP+FP}$
recall The positive precision can be defined as the ratio of correctly predicted positive samples to the total number of samples labeled positive.
$R=\frac{TP}{TP+FN}$
F1-score The F1-score can be defined as the harmonic average of precision and recall.
$F1=\frac{{2}\times{P}\times{R}}{P+R}$
4.4 Experimental Results
The main results of the experiments on clickbait detection in English and Chinese datasets on zero-shot and few-shot scenarios are listed in Table 4, 5 and 6 respectively. We can see that the performance of ChatGPT has not demonstrated the best results on all four metrics over the seven datasets. Compared to the deep neural networks and fine-tuning PLMs, ChatGPT has a lot of room for improvement in clickbait detection. Moreover, with the increasing of pre-trained information, the effectiveness of clickbait detection, sometimes, is worse. Therefore, we can observe that a small amount of pre-training information does not significantly impact the performance of ChatGPT in detecting clickbait. Moreover, we can observe that ChatGPT can achieve stable results on almost all evaluation metrics. The results confirmed that LLMs can be adapted to other languages, and ChatGPT is a monolithic model capable of supporting multiple languages, which makes it a comprehensive multilingual clickbait detection technique.
4.5 Ablation Study
We compared the results between different prompts (as shown in Table 1 and Table 2) for clickbait detection tasks. As same as the main experiments, we select GPT3.5 as a backbone for the prompts experiments. The results comparison of different Prompts is shown in Table 7 and Table 8.
Generally speaking, there is no significant performance gap between the first prompt (P1) and the second prompt (P2) in terms of effectiveness for few-shot clickbait detection. Specifically, the performance of P1 is significantly better than that of P2 in the dataset Wechat. In dataset Sina, the opposite is true.
5 Conclusion and Future Work
In this paper, we present a study of the performance of LLMs (ChatGPT with GPT3.5) for clickbait detection. During the benchmark experiments on both English and Chinese datasets, LLMs did not perform as well as the current state-of-the-art deep neural networks and fine-tuning PLMs methods in the realm of multilingual clickbait detection. In our subsequent efforts, we will try to design more effective methods with the help of LLMs that can significantly and consistently outperform SOTA clickbait detection methods.